If 75% of any increase in income is spent on consumption, then an increase in autonomous investment of R1 billion results in an increase in national income of as much as


R5 billion


R1.33 billion


R4 billion


R6 billion



1
Expert's answer
2021-09-15T11:29:04-0400

The 75% of any increase in income is spent on consumption means the marginal proprnsity to consume(mpc) is 0.75.

Multiplier(k) = 1 / (1 - mpc) = 1 / (1 - 0.75)

Or,  Multiplier(k) = 1 / 0.25 = 4

Now, the increase in autonomous investment (I) is R1 billion.

The increase in national income (Y) will be as follows, Y = [1 / (1 - mpc)] * I

Y = 4 * R1 billion = R4 billion

So, the national income will increase by 4 times of R1 billion or R4 billion.

An increase in autonomous investment of R1 billion results in an increase in national income of R4 billion.